OpenStudy (joshmartin16):

Find the derivative of f(x) = 6 divided by x at x = -2.

zepdrix (zepdrix):

Hey Josh, welcome to OpenStudy! Recall that we have a handy exponent rule to help us out on problems like this, \(\large\rm \dfrac{1}{x}=x^{-1}\) From there we can apply our `power rule`. Have you learned about power rule yet? Or do we need to use the limit definition of the derivative at this point?

OpenStudy (joshmartin16):

no i havent

OpenStudy (joshmartin16):

I want to say the answer is -3/2 but im not sure if its correct

zepdrix (zepdrix):

Ok so we'll have some tricky algebra to take care of then.\[\large\rm f(\color{orangered}{x})=\frac{6}{\color{orangered}{x}}\] If we evaluate this function at x+h, instead of x, it looks like this\[\large\rm f(\color{orangered}{x+h})=\frac{6}{\color{orangered}{x+h}}\]

zepdrix (zepdrix):

\[\large\rm \lim_{h\to0}\frac{f(x+h)-f(x)}{h}\quad=\quad \lim_{h\to0}\frac{\frac{6}{x+h}-\frac{6}{x}}{h}\]

OpenStudy (joshmartin16):

so then its not negative its just 3/2

zepdrix (zepdrix):

I dunno, you're jumping straight to the end lol :)

OpenStudy (joshmartin16):

so what numbers do i input

zepdrix (zepdrix):

\[\large\rm f'(-2)\quad=\quad \lim_{h\to0}\frac{\frac{6}{-2+h}-\frac{6}{-2}}{h}\]I guess -2 is our input. You'll want to do something with the numerator, maybe common denominator, ya?

OpenStudy (joshmartin16):

i dont know how to do that

OpenStudy (salsabyl):

The derivative of 1/x is -1/x squared Multiply by the constant 6 and you havevthe general derivative. Then replace the x by -2

OpenStudy (joshmartin16):

ohhhh okay! give me a second

OpenStudy (joshmartin16):

so (-1/-2) * 6

OpenStudy (salsabyl):

Yeah but the -2 is squared

OpenStudy (joshmartin16):

you cant have exponents on the bottom though?

OpenStudy (salsabyl):

Yes you can If you have exponent -2 for example on top, you can transfer it to the bottom but it becomes 2. The sign of the exponent changes but you can have them both on top or bottom

OpenStudy (joshmartin16):

so the answer is 6!

OpenStudy (salsabyl):

The answer is 6 *(-1/(-2)squared)

OpenStudy (salsabyl):

So -6/4

OpenStudy (salsabyl):

So -3/2

OpenStudy (joshmartin16):

awesome thank you!

OpenStudy (salsabyl):

So your first answer was correct

OpenStudy (salsabyl):

You're welcome :)